A car of 700 kg moving with a velocity of 10 m/s. A motor bike of 120 kg also moving with same speed in the same direction lying behind 10 meter from the car. What is there collision velocity?


1
Expert's answer
2022-03-01T17:33:06-0500

"u=\frac{Mv}{m_1v+m_2v}=10~\frac ms."